# Computing Balanced Solutions for Large International Kidney Exchange Schemes When Cycle Length Is Unbounded

### v1

@@ 3 ILP Formulation
In this section, we show how to find an optimal solution of a partitioned permutation game that is strongly close to a given target allocation $x$ by solving a sequence of Integer Linear Programs (ILPs). Let $(N,v)$ be a partitioned permutation game defined on a directed graph $G=(V,A)$. Recall that for a maximum cycle packing ${\mathcal{C}}$ of $G$, we let $s_{p}({\mathcal{C}})$ denote the number of arcs $(u,v)$ with $v\in V_{p}$ that belong to some directed cycle of ${\mathcal{C}}$. Recall also that $|x_{p}-s_{p}({\mathcal{C}})|$ is the deviation of country $p\in N$ from its target $x_{p}$ if ${\mathcal{C}}$ is chosen as optimal solution. Moreover, in the vector $d({\mathcal{C}})=(|x_{p_{1}}-s_{p_{1}}({\mathcal{C}})|,\dots,|x_{p_{n}}-s_{p_{n}}({\mathcal{C}})|)$, the deviations $|x_{p}-s_{p}({\mathcal{C}})|$ are ordered non-increasingly. Finally, we recall that ${\mathcal{C}}$ is strongly close to $x$ if $d({\mathcal{C}})$ is lexicographically minimal over all optimal solutions for $(N,v)$.
In the kidney exchange literature the following ILP is called the edge-formulation (see e.g. [1]) For each $a\in A$, let $e_{a}\in\{0,1\}$ be a binary edge-variable for $D$.
@@
This ILP has $|A|$ binary variables and $2|V|$ constraints. In the following we are going to sequentially find largest country deviations $d_{t}^{*}$ ($t\geq 1$) and the corresponding minimal number $n_{t}^{*}$ of countries receiving that deviation. We achieve this by solving one ILP of similar size for each $d_{t}^{*}$ and $n_{t}^{*}$, so two ILPs per iteration $t$. By similar size, we mean that in each iteration we are going to add $|N|$ binary variables and a single additional constraint, while $|N|\leq|V|$ holds by definition and typically $|N|$ is much smaller than $|A|$. Meanwhile, since at every iteration we are going to fix the deviation of at least one additional country (we will not necessarily know which country, we are only going to keep track of number of countries with fixed deviation), the number of iterations are at most $|N|$ (as $t\leq|N|$). Hence, we will solve no more than $2|N|$ ILPs, among which the largest has $\mathcal{O}(|A|+|N|^{2})$ binary variables and $\mathcal{O}(|V|+|N|)=\mathcal{O}(|V|)$ constraints.
@@
The first three constraints guarantee that all solutions are in fact maximum cycle-packings. These constraints will be part of the formulation throughout the entire ILP-series. The remaining two constraints, together with the objective function guarantees that we minimize the largest country deviation. Note that for each country $p$, exactly one of $\sum_{j\in V_{p}}e_{ij}-x_{p}$ and $x_{p}-\sum_{j\in V_{p}}e_{ij}$ is positive and exactly one of them is negative, unless both of them are zero. However, as soon as we reach $d_{t}^{*}\leq 1/2$ we have found a strongly close maximum cycle-packing. Hence, in the remainder, we assume $d_{t}^{*}>1/2$ for each $t$ such that the series continues with $t+1$.
($\text{ILP}_{d_{1}}$) has one additional continuous variable ($d_{1}$) and $12|N|+$ 1 additional constraints. For every country $p\in N$ we have that $|\sum_{j\in V_{p}}e_{ij}-x_{p}|\leq d_{1}^{*}$. However, there exists a smallest subset $N_{1}\subseteq N$ (which may not necessarily be unique) such that
@@
In a solution of ($\text{ILP}_{d_{1}}$), let $n_{1}$ be the number of countries with deviation $d_{1}^{*}$. We need to determine if there is another solution of ($\text{ILP}_{d_{1}}$) with fewer than $n_{1}$, possibly $n_{1}^{*}$, countries having $d_{1}^{*}$ deviation. For this purpose we must be able to distinguish between countries unable to have less than $d_{1}^{*}$ deviation and countries for which the deviation is at most $d_{2}^{*}$, the latter value unknown at this stage. In order to make this distinction, we determine a lower bound on $d_{1}^{*}-d_{2}^{*}$ by examining the target allocation $x$. We will then set $\varepsilon$ in the next ILP to be strictly smaller than this lower bound.
The number of vertices for a country covered by any cycle-packing is an integer. Hence, the number of possible country deviations is at most $2|N|$ and depends only on $x$. The fractional part of a country deviation is either $\operatorname{frac}(x_{p})$ or $1-\operatorname{frac}(x_{p})$. Therefore, to find the minimal positive difference in between the deviations of any two countries $p$ and $q$, we have to compare the values $\operatorname{frac}(x_{p})$ and $1-\operatorname{frac}(x_{p})$, with $\operatorname{frac}(x_{q})$ or $1-\operatorname{frac}(x_{q})$ and take the minimum of those four possible differences. Let $\varepsilon$ be a small positive constant that is strictly smaller than the minimum possible positive difference between any two countries.
We will distinguish between countries having minimal deviation of $d_{1}^{*}$ and others through additional binary variables. Since later in the ILP series we will need to distinguish between countries fixed at different deviation levels, let us introduce $z^{t}_{p}\in\{0,1\}$ binary variables, where $z^{t}_{p}=1$ indicates that $p\in N_{t}$.
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\ \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\ \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}-\varepsilon(1-z^{1}_{p})&\forall p\in N\\ \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}-\varepsilon(1-z^{1}_{p})&\forall p\in N\end{array}$
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\ \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\ \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}&\forall p\in N\\ \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}&\forall p\in N\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{2}+z^{1}_{p}d_{1}^{*}&\forall p\in N\\ \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{2}+z^{1}_{p}d_{1}^{*}&\forall p\in N\\ \displaystyle\sum_{p\in N}z^{1}_{p}&=&n_{1}^{*}&\end{array}$
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\ \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\ \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{2}^{*}-\varepsilon(1-z^{2}_{p})+z_{p}^{1}L&\forall p\in N\\ \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{2}^{*}-\varepsilon(1-z^{2}_{p})+z_{p}^{1}L&\forall p\in N\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}&\forall p\in N\\ \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}&\forall p\in N\\ \displaystyle\sum_{p\in N}z^{1}_{p}&=&n_{1}^{*}&\\ z_{p}^{1}+z_{p}^{2}&\leq&1&\forall p\in N\end{array}$
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\ \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\ \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\ \displaystyle\sum_{i=1}^{t-1}z^{i}_{p}&\leq&1&\forall p\in N\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{t}+\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}d_{i}^{*}&\forall p\in N\\ \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{t}+\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}d_{i}^{*}&\forall p\in N\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}d_{i}^{*}+(1-\displaystyle\sum_{i=1}^{t-1}z^{i}_{p})L&\forall p\in N\\ \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&\displaystyle\sum_{i=1}^{t-1}z^{i}_{p}d_{i}^{*}+(1-\displaystyle\sum_{i=1}^{t-1}z^{i}_{p})L&\forall p\in N\\ \displaystyle\sum_{p\in N}z^{i}_{p}&=&n_{i}^{*}&\forall i=1,\dots,t-1\end{array}$
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\ \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\ \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\ \displaystyle\sum_{i=1}^{t}z^{i}_{p}&\leq&1&\forall p\in N\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{t}^{*}-\varepsilon(1-z^{t}_{p})+\displaystyle\sum_{i=1}^{t-1}z_{p}^{i}d_{i}^{*}&\forall p\in N\\ x_{p}-\displaystyle\sum_{j\in V_{p}}e_{ij}&\leq&d_{t}^{*}-\varepsilon(1-z^{t}_{p})+\displaystyle\sum_{i=1}^{t-1}z_{p}^{i}d_{i}^{*}&\forall p\in N\\ \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&\displaystyle\sum_{i=1}^{t}z_{p}^{i}d_{i}^{*}+(1-\displaystyle\sum_{i=1}^{t}z_{p}^{i})L&\forall p\in N\\ \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&\displaystyle\sum_{i=1}^{t}z_{p}^{i}d_{i}^{*}+(1-\displaystyle\sum_{i=1}^{t}z_{p}^{i})L&\forall p\in N\\ \displaystyle\sum_{p\in N}z^{i}_{p}&=&n_{i}^{*}&\forall i=1,\dots,t-1\end{array}$
@@ 4 Simulations
    to examine the benefits of strongly close optimal solutions over weakly close optimal solutions or arbitrary chosen optimal solutions when $\ell=\infty$;
@@
    to examine benefits of using credits when $\ell=\infty$;
@@
    to examine the exchange cycle length distribution when $\ell=\infty$; and
@@
    to compare the results for $\ell=\infty$ with the known results [8] for the other extreme case when $\ell=2$.
For our fourth aim, we want to know in particular how much scope there is for improvement in the total number of kidneys when we move from $\ell=2$ to $\ell=\infty$.
@@
Set Up. To do a fair comparison we follow the same set up as in [8], and in addition, we also use the data from [6] that was used in [8]. That is, for our simulations, we take the same 100 compatibility graphs $G_{1},\ldots,G_{100}$, each with roughly 2000 vertices from [6, 8]. As real medical data is unavailable to us for privacy reasons, the data from [6, 8] was obtained by using the data generator from [33]. This data generator was used in many papers and is the most realistic synthetic data generator available; see also [23].
@@
For every $i\in\{1,\ldots,100\}$ we do as follows. For every $n\in\{4,\ldots,10\}$, we perform simulations for $n$ countries. We first partition $V(G_{i})$ into the same $n$ sets $V_{i,1},\ldots,V_{i,n}$ as in [8] of equal size ${2000}/{n}$ (subject to rounding), so $V_{i,p}$ is the set of patient-donor pairs of country $p$. For round 1, we construct a compatibility graph $G_{i}^{1}$ as a subgraph of $G_{i}$ of size roughly $500$. We add the remaining patient-donor pairs of $G_{i}$ as vertices by a uniform distribution between the remaining rounds.
Starting with $G_{i}^{1}(n)$ we run a IKEP of 24 rounds in total. This gives us 24 compatibility graphs $G_{i}^{1}(n),\ldots,G_{i}^{24}(n)$. Just as in [8], any patient-donor pair, whose patient is not helped within four rounds, will be automatically deleted from the pool.
@@
A (24-round) simulation instance consists of the data needed to generate a graph $G_{i}^{1}(n)$ and its successors $G_{i}^{2}(n),\ldots,G_{i}^{24}(n)$, together with specifications for the choice of initial allocation and optimal solution (maximum cycle packing). Our code for obtaining the simulation instances is in GitHub repository [42], along with the compatibility graphs data and the seeds for the randomization.
We now discuss our choice for the initial allocations and optimal solutions (maximum cycle packings).

### v2

@@ 3 ILP Formulation
In this section, we show how to find an optimal solution of a partitioned permutation game that is strongly close to a given target allocation $x$ by solving a sequence of Integer Linear Programs (ILPs). Let $(N,v)$ be a partitioned permutation game defined on a directed graph $G=(V,A)$ with vertex partition $V_{1},\ldots,V_{n}$ for some $n\geq 1$. Recall that for a maximum cycle packing ${\mathcal{C}}$ of $G$, we let $s_{p}({\mathcal{C}})$ denote the number of arcs $(u,v)$ with $v\in V_{p}$ that belong to some directed cycle of ${\mathcal{C}}$. Recall also that $|x_{p}-s_{p}({\mathcal{C}})|$ is the deviation of country $p\in N$ from its target $x_{p}$ if ${\mathcal{C}}$ is chosen as optimal solution. Moreover, in the vector $d({\mathcal{C}})=(|x_{p_{1}}-s_{p_{1}}({\mathcal{C}})|,\dots,|x_{p_{n}}-s_{p_{n}}({\mathcal{C}})|)$, the deviations $|x_{p}-s_{p}({\mathcal{C}})|$ are ordered non-increasingly. Finally, we recall that ${\mathcal{C}}$ is strongly close to $x$ if $d({\mathcal{C}})$ is lexicographically minimal over all optimal solutions for $(N,v)$.
In the kidney exchange literature the following ILP is called the edge-formulation (see e.g. [1]). For each $a\in A$ in the graph $D$, let $e_{a}\in\{0,1\}$ be a (binary) edge-variable; this yields a vector $e$ with entries $e_{a}$.
@@
This ILP has $|A|$ binary variables and $2|V|$ constraints. In the following we are going to sequentially find largest country deviations $d_{t}^{*}$ ($t\geq 1$) and the corresponding minimal number $n_{t}^{*}$ of countries receiving that deviation. We achieve this by solving an ILP of similar size for each $d_{t}^{*}$ and $n_{t}^{*}$, so two ILPs per iteration $t$. By similar size, we mean that in each iteration we are going to add $|N|$ binary variables and a single additional constraint, while $|N|\leq|V|$ holds by definition and typically $|N|$ is much smaller than $|A|$. Meanwhile, since at every iteration we are going to fix the deviation of at least one additional country (we will not necessarily know which country, we are only going to keep track of the number of countries with fixed deviation), the number of iterations are at most $|N|$ (as $t\leq|N|$). Hence, we will solve no more than $2|N|$ ILPs, among which the largest has $O(|A|+|N|^{2})$ binary variables and $O(|V|+|N|)=O(|V|)$ constraints.
@@
The first three constraints guarantee that all solutions are, in fact, maximum cycle packings. These constraints will be part of the formulation throughout the entire ILP-series. The remaining two constraints, together with the objective function, guarantee that we minimize the largest country deviation. Note that for each country $p$, exactly one of $\sum_{j\in V_{p}}e_{ij}-x_{p}$ and $x_{p}-\sum_{j\in V_{p}}e_{ij}$ is positive and exactly one is negative, unless both are zero. However, as soon as we reach $d_{t}^{*}\leq 1/2$ we have found a strongly close maximum cycle packing. Hence, in the remainder, we assume $d_{t}^{*}>1/2$ for each $t$ such that the series continues with $t+1$.
($\text{ILP}_{d_{1}}$) has one additional continuous variable ($d_{1}$) and $12|N|+$ 1 additional constraints. For every country $p\in N$, we have that $|\sum_{j\in V_{p}}e_{ij}-x_{p}|\leq d_{1}^{*}$. However, there exists a smallest subset $N_{1}\subseteq N$ (which may not necessarily be unique) such that
@@
In a solution of ($\text{ILP}_{d_{1}}$), let $n_{1}$ be the number of countries with deviation $d_{1}^{*}$. We need to determine if there is another solution of ($\text{ILP}_{d_{1}}$) with fewer than $n_{1}$, possibly $n_{1}^{*}$, countries having deviation $d_{1}^{*}$. For this purpose we must be able to distinguish between countries unable to have less than $d_{1}^{*}$ deviation and countries for which the deviation is at most $d_{2}^{*}$, the latter value unknown at this stage. In order to make this distinction, we determine a lower bound on $d_{1}^{*}-d_{2}^{*}$ by examining the target allocation $x$. We will then set $\varepsilon$ in the next ILP to be strictly smaller than this lower bound.
The number of vertices for a country covered by any cycle packing is an integer. Hence, the number of possible country deviations is at most $2|N|$ and depends only on $x$. The fractional part of a country deviation is either $\operatorname{frac}(x_{p})$ or $1-\operatorname{frac}(x_{p})$. Therefore, to find the minimal positive difference in between the deviations of any two countries $p$ and $q$, we have to compare the values $\operatorname{frac}(x_{p})$ and $1-\operatorname{frac}(x_{p})$, with $\operatorname{frac}(x_{q})$ or $1-\operatorname{frac}(x_{q})$ and take the minimum of those four possible differences. Let $\varepsilon$ be a small positive constant that is strictly smaller than the minimum possible positive difference between any two countries.
We will distinguish between countries having minimal deviation of $d_{1}^{*}$ and others through additional binary variables. Since later in the ILP series we will need to distinguish between countries fixed at different deviation levels, let us introduce binary variables $z^{t}_{p}$, where $z^{t}_{p}=1$ indicates that $p\in N_{t}$.
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[15.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[15.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}-\varepsilon(1-z^{1}_{p})&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}-\varepsilon(1-z^{1}_{p})&\forall p\in N\end{array}$
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[15.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[15.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{2}+z^{1}_{p}d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{2}+z^{1}_{p}d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle\sum_{p\in N}z^{1}_{p}&=&n_{1}^{*}&\end{array}$
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[15.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[15.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{2}^{*}-\varepsilon(1-z^{2}_{p})+z_{p}^{1}L&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{2}^{*}-\varepsilon(1-z^{2}_{p})+z_{p}^{1}L&\forall p\in N\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&d_{1}^{*}&\forall p\in N\\[15.0pt] \displaystyle\sum_{p\in N}z^{1}_{p}&=&n_{1}^{*}&\\[15.0pt] z_{p}^{1}+z_{p}^{2}&\leq&1&\forall p\in N\end{array}$
@@
[…]
@@
$\begin{array}[]{rclr}\displaystyle\sum_{j:ji\in A}e_{ji}&=&\displaystyle\sum_{j:ij\in A}e_{ij}&\forall i\in V\\[15.0pt] \displaystyle\sum_{j:ji\in A}e_{ji}&\leq&1&\forall i\in V\\[15.0pt] \displaystyle\sum_{ij\in A}e_{ij}&=&M^{*}&\\[15.0pt] \displaystyle\sum_{i=1}^{t}z^{i}_{p}&\leq&1&\forall p\in N\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&d_{t}^{*}-\varepsilon(1-z^{t}_{p})+\displaystyle\sum_{i=1}^{t-1}z_{p}^{i}d_{i}^{*}&\forall p\in N\\[15.0pt] x_{p}-\displaystyle\sum_{j\in V_{p}}e_{ij}&\leq&d_{t}^{*}-\varepsilon(1-z^{t}_{p})+\displaystyle\sum_{i=1}^{t-1}z_{p}^{i}d_{i}^{*}&\forall p\in N\\[15.0pt] \displaystyle\sum_{j\in V_{p}}e_{ij}-x_{p}&\leq&\displaystyle\sum_{i=1}^{t}z_{p}^{i}d_{i}^{*}+(1-\displaystyle\sum_{i=1}^{t}z_{p}^{i})L&\forall p\in N\\[15.0pt] \displaystyle x_{p}-\sum_{j\in V_{p}}e_{ij}&\leq&\displaystyle\sum_{i=1}^{t}z_{p}^{i}d_{i}^{*}+(1-\displaystyle\sum_{i=1}^{t}z_{p}^{i})L&\forall p\in N\\[20.0pt] \displaystyle\sum_{p\in N}z^{i}_{p}&=&n_{i}^{*}&\forall i=1,\dots,t-1\end{array}$
@@ 4 Simulations
    to examine the benefits of strongly close optimal solutions over weakly close optimal solutions or arbitrarily chosen optimal solutions;
@@
    to examine the benefits of using credits;
@@
    to examine the exchange cycle length distribution; and
@@
    to compare all these results with those for the other extreme case where $\ell=2$ [10].
For our fourth aim, we are especially interested in the increase in the total number of patients treated when we compare the $\ell=2$ and $\ell=\infty$ cases.
4.1 The Set Up
@@
To allow a fair comparison we use the same set up as in [10], and in addition, we also use the data from [6] that was used in [10]. That is, for our simulations, we take the same 100 compatibility graphs $G_{1},\ldots,G_{100}$, each with roughly 2000 vertices from [6, 10]. As real medical data is unavailable to us for privacy reasons, the data from [6] was obtained using the data generator from [34]. This data generator was used in many papers and is the most realistic synthetic data generator available; see also [24].
@@
We use every $G_{i}$ ($i\in\{1,\ldots,100\}$) as the basis to perform simulations for $n$ countries, where we let $n=4,\ldots,10$. In order to do this, we first consider the case where all countries have equal size. That is, we partition $V(G_{i})$ into the same $n$ sets $V_{i,1},\ldots,V_{i,n}$ as in [10] which are of equal size ${2000}/{n}$ (subject to rounding), so $V_{i,p}$ is the set of patient-donor pairs of country $p$. For round 1, we construct a compatibility graph $G_{i}^{1}(n)$ as a subgraph of $G_{i}$ of size roughly $500$. We add the remaining patient-donor pairs of $G_{i}$ as vertices by a uniform distribution between the remaining rounds $2,\ldots,24$. So, starting with $G_{i}^{1}(n)$, we run an IKEP of 24 rounds in total. This gives us 24 compatibility graphs $G_{i}^{1}(n),\ldots,G_{i}^{24}(n)$. As in [10], any patient-donor pair whose patient is not helped within four rounds will automatically be deleted from the pool.
We also consider a situation where countries have different sizes. We note that there are numerous ways one could allow for differently sized countries. However, we follow the approach of [10], as this allows us to compare the results with those for $\ell=2$. That is, we consider the same 100 compatibility graphs $D_{1},\ldots,D_{100}$ as before; the only difference is that we now partition each $V(D_{i})$ into the same $n$ sets $V_{i,1},\ldots,V_{i,n}$ as in [10], such that
  • approximately $n/3$ are small, that is, have size roughly $1000/n$ (subject to rounding);
  • approximately $n/3$ are medium, that is, have size roughly $2000/n$; and
  • approximately $n/3$ are large, that is, have size roughly $3000/n$.
A (24-round) simulation instance consists of
  • (i)
    the data needed to generate a graph $G_{i}^{1}(n)$ and its successors $G_{i}^{2}(n),\ldots,G_{i}^{24}(n)$
  • (ii)
    an indication of whether the countries must be of the same size or can have different sizes (in the way explained above);
  • (iii)
    a choice of solution concept for computing the initial allocation (see Section 4.2); and
  • (iv)
    a choice of type of optimal solution (see Section 4.3).
@@
Our code for obtaining the simulation instances can be found in a GitHub repository [44], along with the compatibility graphs data and the seeds for the randomization.
We now discuss our choice for the initial allocations and optimal solutions.
4.2 The Initial Allocations
For the initial allocations $y$ we use seven known solution concepts, which we all define below:
  • the Shapley value,
  • the Banzhaf value,
  • the nucleolus,
  • the tau value,
  • the benefit value
  • the contribution value, and
  • the Banzhaf* value (which, we recall, is a conceptually different solution concept).
We define each of these solution concepts below. Each of them prescribes exactly one allocation for the partitioned permutation games associated with the computability graphs. For computing these allocations we need the $v$-values of the partitioned permutation games.
We can compute a single value $v(S)$ in polynomial time by Theorem 1.1 through a transformation into a bipartite graph and solving a maximum weight perfect matching problem (see Section 1). For this we use the package LEMON version 1.3.1 of [25].
Let $(N,v)$ be a cooperative game. The Shapley value $\phi(N,v)$ [37] is defined by setting for $p\in N$,
$\phi_{p}(N,v)=\sum_{S\subseteq N\backslash\{p\}}\frac{|S|!(n-|S|-1)!}{n!}(v(S\cup\{p\})-v(S)).$
To define the next solution concept, we first introduce unnormalized Banzhaf value $\psi(N,v)$ [31] defined by setting for $p\in N$,
$\psi_{p}(N,v):=\sum_{S\subseteq N\backslash\{p\}}\frac{1}{2^{n-1}}(v(S\cup\{p\})-v(S)).$
As $\psi_{p}$ may not be an allocation, the (normalized) Banzhaf value $\overline{\psi}(N,v)$ of a game $(N,v)$ was introduced and defined by setting for $p\in N$,
$\overline{\psi}_{p}(N,v):=\frac{\psi_{p}(N,v)}{\sum_{q\in N}\psi_{q}(N,v)}\cdot v(N).$